OpenStudy (anonymous):

Find f inverse if f(x) = x^3 + 2 PLease HELP!!

OpenStudy (anonymous):

This can be seen as: \[y = x^3 + 2\] To find the inverse you must swap x for y and solve for y. \[x = y^3 + 2\]

OpenStudy (anonymous):

@waterineyes was yours not correct?

OpenStudy (anonymous):

It was correct..

OpenStudy (anonymous):

His was 100% correct. It's important to see why.

OpenStudy (anonymous):

why did you delete it?

OpenStudy (anonymous):

I again provide you the solution: Now I am starting where @qpHalcy0n left: \[x = y^3 + 2\] \[y^3 = x-2\] \[\large \color{green}{y = \sqrt[3]{x - 2}}\]

OpenStudy (anonymous):

Now replace y by \(f^{-1}(x)\).

OpenStudy (anonymous):

so final answer is f^-1x=cubed root x-2? @waterineyes

OpenStudy (anonymous):

Yes it is the answer..

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Welcome dear..